County officials are interested in dog breeders in the county. They collect census data and create a probability distribution.

Number of Dog Types P(X)
1 0.55
2 0.2
3 0
4 0.2
5 0.05

Find the expected number of dog-types in the county. Answer with a simplified fraction.
Answer #1

Solution :

GIVEN TABLE

Number of Dog Types P(X)
1 0.55
2 0.2
3 0
4 0.2
5 0.05

expected value = E(X) = X * P(X)

= 1 *0.55 + 2 *0.2 + 3 *0 + 4 *0.2 +5*0.05

= (0.55 +0.4 +0 + 0.8+0.25)

= 2